OpenStudy (coolguy34):

i already have the first step done i just need the rest of the steps: 12x − 2(y− 1) = 10

OpenStudy (alexandervonhumboldt2):

what do you need to solve for?

OpenStudy (coolguy34):

y

OpenStudy (coolguy34):

is what i need to solve for

OpenStudy (alexandervonhumboldt2):

12x − 2(y− 1) = 10 use PEDMAS first distribute: 12x-2y+2=10 now subtract 2 from both sides: 12x-2y+2-2=10-2 12x-2y=8 now subtract 12x from both sides: 12x-12x-2y=8-12x -2y=8-12x divide both sides by -2: -2y/-2=(8-12x)/-2 y=-4+6x

OpenStudy (coolguy34):

thanks